# SoulX-Singer: Towards High-Quality Zero-Shot Singing Voice Synthesis

### v1

@@ Abstract
While recent years have witnessed rapid progress in speech synthesis, open-source singing voice synthesis (SVS) systems still face significant barriers to industrial deployment, particularly in terms of robustness and zero-shot generalization. In this report, we introduce SoulX-Singer, a high-quality open-source SVS system designed with practical deployment considerations in mind. SoulX-Singer supports controllable singing generation conditioned on either symbolic musical scores (MIDI) or melodic representations, enabling flexible and expressive control in real-world production workflows. Trained on more than 42,000 hours of vocal data, the system supports Mandarin Chinese, English, and Cantonese, and consistently achieves state-of-the-art synthesis quality across languages under diverse musical conditions. Furthermore, to enable reliable evaluation of zero-shot SVS performance in practical scenarios, we construct SoulX-Singer-Eval, a dedicated benchmark with strict training–test disentanglement, facilitating systematic assessment in zero-shot settings.

### v2

@@ Abstract
While recent years have witnessed rapid progress in speech synthesis, open-source singing voice synthesis (SVS) systems still face significant barriers to industrial deployment, particularly in terms of robustness and zero-shot generalization. In this report, we introduce SoulX-Singer, a high-quality open-source SVS system designed with practical deployment considerations in mind. SoulX-Singer supports controllable singing generation conditioned on either symbolic musical scores (MIDI) or melodic representations, enabling flexible and expressive control in real-world production workflows. Trained on more than 42,000 hours of vocal data, the system supports Mandarin Chinese, English, and Cantonese and consistently achieves state-of-the-art synthesis quality across languages under diverse musical conditions. Furthermore, to enable reliable evaluation of zero-shot SVS performance in practical scenarios, we construct SoulX-Singer-Eval, a dedicated benchmark with strict training-test disentanglement, facilitating systematic assessment in zero-shot settings.
@@ 6 Acknowledgments
Appendix
SoulX-Singer-SVC
To extend SoulX-Singer to scenarios where precise MIDI and time-aligned lyrics are unavailable, we further introduce SoulX-Singer-SVC, a singing voice conversion (SVC) model derived from SoulX-Singer. Specifically, we replace the original score and text encoders with a frozen Whisper-base encoder to extract robust semantic representations directly from the source singing audio. Meanwhile, the F0 encoder and the flow-matching decoder are initialized with pre-trained SoulX-Singer parameters to preserve melody controllability and high-fidelity synthesis quality. The model is then fine-tuned on a relatively small but high-quality singing voice dataset.
We evaluate SoulX-Singer-SVC on GMO-SVS and SoulX-Singer-Eval, comparing it with GroundTruth, melody-controlled SoulX-Singer, and state-of-the-art SVC systems, including YingMusic-SVC [2] and Vevo-SVC [27]. Note that the FFE metric is omitted for SoulX-Singer-Eval because the source audio was pitch-shifted. The results are presented in Table 4, and the model is released as part of the SoulX-Singer project.
Table 4: Singing voice conversion performance results on two datasets ($\uparrow$ larger is better, $\downarrow$ smaller is better). The best and the second best results are shown in bold and underlined.
GMO-SVS Dataset
Model Chinese English
[…]
GroundTruth 0.074 - - 4.624 4.334 0.197 - - 4.441 3.825
SoulX-Singer (Melody) 0.065 0.897 0.044 4.458 4.110 0.151 0.918 0.036 4.323 3.751
YingMusic-SVC 0.097 0.908 0.057 4.399 3.974 0.243 0.923 0.046 4.337 3.708
Vevo-SVC 0.143 0.892 0.104 4.471 4.088 0.334 0.920 0.079 4.389 3.748
SoulX-Singer-SVC 0.107 0.904 0.036 4.419 4.029 0.267 0.919 0.030 4.314 3.666
SoulX-Singer-Eval Dataset
Model Chinese English
[…]
GroundTruth 0.089 - - 4.450 4.237 0.208 - - 4.569 3.815
SoulX-Singer (Melody) 0.072 0.869 - 4.402 4.061 0.161 0.883 - 4.249 3.755
YingMusic-SVC 0.096 0.876 - 4.227 3.848 0.236 0.881 - 4.274 3.727
Vevo-SVC 0.180 0.871 - 4.387 3.932 0.346 0.886 - 4.359 3.772
SoulX-Singer-SVC 0.102 0.880 - 4.303 3.938 0.260 0.894 - 4.286 3.774